# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

### v1

@@ 3 Reduced Order Hysteretic Magnetization Model
This model is adapted and inspired from the energy-based model for ferromagnetic materials [27, 28].
We start in Section 3.1 by clarifying how the ROHM model can be used to model the strand. We then define the building blocks, referred to as cells, that will be combined for the construction of the model. Each cell defines a local hysteretic relationship between the magnetic field $\boldsymbol{h}$ and the magnetic flux density $\boldsymbol{b}$. We present in Section 3.2 a rate-independent cell, that we call the superconductor cell (S cell). In Section 3.3, we add contributions to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell). In Section 3.4, we discuss how to reproduce the effect of a field-dependent critical current density within the cells. Finally, we show in Section 3.5 how several cells can be combined into a chains of cell, defining a complete ROHM model.
3.1 Concept of a reduced order model
[…]
@@ 3.1 Concept of a reduced order model
Similarly, the power loss is no longer obtained via a current density as in Eq. (2), but is now instead described by the magnetic power density $\boldsymbol{h}\cdot\dot{\boldsymbol{b}}$. Part of this power is related to stored magnetic energy, and hence not associated with loss, and the remaining part is related to dissipated energy. One major advantage of the energy-based hysteresis model used in this work is that it gives access to both parts separately at any time [27].
@@
where $\boldsymbol{h}_{\text{in}}$ has no immediate meaning in the detailed strand model per se, but is the local magnetic field the hysteresis model will see inside the strand. These curves are shown in Fig. 7. At high frequencies ($f\gtrsim 10^{4}$ Hz), they approach the shape of an ellipse with a slope of $-1$. The area of closed loops are identical in both Figs. 5 and 7, as shown in B.
Figure 7: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T. Same results as in Fig. 5 but in terms of $\mu_{0}h_{\text{in}}$ instead of $\mu_{0}h_{\text{app}}$ (see Eq. (7)).
@@ 3.2 Superconductor cell - S cell
As a first step, we present a rate-independent hysteresis cell [27], that we refer to as the superconductor cell, or S cell. In this cell, the magnetic field $\boldsymbol{h}$ is decomposed into a reversible field, $\boldsymbol{h}_{\text{rev}}$, and an irreversible field, $\boldsymbol{h}_{\text{irr}}$:
@@
The reversible field defines the magnetic flux density:
@@
The irreversible field creates hysteresis by introducing history dependence. Its amplitude $\|\boldsymbol{h}_{\text{irr}}\|$ is bounded by a value $\kappa$ (A/m), the irreversibility parameter. From a known magnetic field $\boldsymbol{h}$, which is the driving vector, and a given reversible field $\boldsymbol{h}_{\text{rev}}$ that has been established previously (as a function of the history of $\boldsymbol{h}$), the irreversible field $\boldsymbol{h}_{\text{irr}}$ is determined as follows:
@@
where the dot notation (as in $\dot{\boldsymbol{b}}$) represents the time derivative of the quantity. An illustration of this equation is given in Fig. 8. While the driving field $\boldsymbol{h}$ evolves inside the sphere of radius $\kappa$ centered at $\boldsymbol{h}_{\text{rev}}$, the reversible field $\boldsymbol{h}_{\text{rev}}$ stays constant, and the irreversible field is defined accordingly, see Fig. 8(a). If $\boldsymbol{h}$ reaches the surface of the sphere, then, the reversible field $\boldsymbol{h}_{\text{rev}}$ must evolve as well, in order to maintain the condition $\|\boldsymbol{h}_{\text{irr}}\|\leq\kappa$ valid, and $\boldsymbol{h}_{\text{irr}}$ is established parallel to $\dot{\boldsymbol{b}}$, see Fig. 8(b).
@@
Figure 8: S cell, adapted from [27, 29]. Illustration of the magnetic field decomposition into reversible and irreversible fields in 2D. The dotted elements in (b) refer to the situation at a previous time.
The total power density $p_{\text{tot}}$ (in W/m3) is expressed as
@@
The reversible power $p_{\text{rev}}$ is the time derivative of a stored magnetic energy density,
@@
while the irreversible power $p_{\text{irr}}$ is always non-negative and is associated with dissipated energy density,
@@
A mechanical analogy of a S cell in a one-dimensional setting is shown in Fig. 9. It represents a mechanical system made of a dry friction element in parallel with a restoring spring [27]. Examples of hysteresis curves computed with the S cell are shown in Fig. 10. These curves are independent of the rate of change of the applied field.
Figure 9: Mechanical analogy [27] of the S cell: a parallel connection of a dry friction element and a restoring spring. The system is subject to a force $h$, which causes an elongation $b=\mu_{0}h_{\text{rev}}$.
Figure 10: Hysteresis curves for the S cell subject to a unidirectional oscillating magnetic field $h$ as an input. The legend is the same for both subfigures. (Left) Magnetic flux density $b$ versus applied field $\mu_{0}h$. (Right) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$.
@@ 3.3 Composite superconductor cell - CS cell
In regimes in which the superconducting strand response is clearly not rate-independent, the S cell presented in the previous section must be generalized. Starting from the S cell, we propose to account for the magnetization due to eddy currents and coupling currents by adding two new contributions in the magnetic field decomposition, $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$, such that Eq. (8) is rewritten
@@
Adapting [28], the field $\boldsymbol{h}_{\text{eddy}}$ is defined as
with an eddy current time constant parameter $\tau_{\text{e}}$ (s). This component produces a loss contribution, $p_{\text{eddy}}$, which is non-negative. It reads:
In the mechanical analogy, the inclusion of eddy current effects corresponds to adding a dashpot (viscous friction element) in parallel with the dry friction element and the restoring spring, as illustrated in Fig. 11(a). For slow variations of the force $h$, the dashpot is barely opposing the elongation $b$, whereas for faster variations of $h$, it becomes gradually more blocking and starts to oppose to variations of $b$. This reproduces magnetic shielding and increased magnetization due to eddy currents.
The next new contribution $\boldsymbol{h}_{\text{coupling}}$ is associated with coupling currents [5]. The coupling currents in the conducting matrix are fed by currents flowing in the superconducting filaments, and are therefore also subject to hysteresis and saturation of the filaments. The field $\boldsymbol{h}_{\text{coupling}}$ has then its own irreversibility parameter $\chi$ (in A/m). We define
$\boldsymbol{h}_{\text{coupling}}$ =$&\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}},&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|<\chi,\\ &\chi\ \dot{\boldsymbol{h}}_{\text{rev}}/\|\dot{\boldsymbol{h}}_{\text{rev}}\|,&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|\geq\chi.\\ \end{aligned}\right.$ (18)
$=\frac{\tau_{\text{c}}}{\mu_{0}}\dot{\boldsymbol{b}}_{\text{coupling}},$ (19)
with a coupling currents time constant $\tau_{\text{c}}$ (s). The field $\boldsymbol{h}_{\text{coupling}}$ introduces two new loss contributions, $p_{\text{coupling}}$ and $p_{\text{irr,c}}$, that read:
$p_{\text{irr,c}}$ $=\boldsymbol{h}_{\text{coupling}}\cdot\left(\dot{\boldsymbol{b}}-\dot{\boldsymbol{b}}_{\text{coupling}}\right).$ (21)
The mechanical analogy associated with this new problem is represented in Fig. 11(a). The combination of the dashpot in series with the dry friction element reproduces the saturation of the coupling currents. When the force on the dashpot remains below $\chi$, the dry friction element is at rest and produces no hysteresis. This corresponds to a situation in which the coupling currents are small enough for the hysteresis to be neglected. If the force exceeds $\chi$ on the dashpot, it is then subject to hysteresis. The second hysteresis element, with $\chi$, always acts in parallel with the first one, related to filament magnetization. The magnetization of coupled filaments therefore depends of the sum of both.
@@
Figure 11: CS cell accounting for eddy currents and coupling current effects. The dashed circle of radius $\chi$ limits the amplitude of the field $\boldsymbol{h}_{\text{coupling}}$.
@@
An illustration of magnetization curves obtained with a CS cell is given in Fig. 12, and the loss per cycle as a function of frequency is given in Fig. 13. The loss per cycle curve exhibits the sought features for the composite strand. The filament curve shows the expected evolution, two plateaus at low and middle frequencies, with a transition where the coupling loss peaks, and then a sharp decrease when the eddy loss becomes dominant. Coupling and eddy loss curves both follow the shape of two bell curves.
Figure 12: Hysteresis curves for the composite hysteresis cell reproducing eddy and coupling currents, with $\kappa=0.2/\mu_{0}$ A/m, $\chi=0.3/\mu_{0}$ A/m, $\tau_{\text{c}}=0.1$ s, and $\tau_{\text{e}}=70$ $\upmu$s. Unidirectional oscillating magnetic field $h$ as an input. The legend is shared between both subfigures. (Left) Magnetic flux density $b$ versus applied field $\mu_{0}h$. (Right) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$.
Figure 13: Total loss per cycle and separated loss contributions as a function of the frequency $f$, for a magnetic field amplitude of $2$ T obtained with the same composite hysteresis cell as in Fig. 12.
Note that the eddy current curve for the eddy loss per cycle should be asymptotically decreasing as $1/\sqrt{f}$ at high frequencies, due to the skin effect. With the CS cell, one can however only expect the curve to decrease as $1/f$ at high frequencies. For $f\to\infty$, the CS cell is therefore not expected to provide perfectly valid results. It should be further improved if a more realistic frequency response is wanted in such high frequencies. For most practical applications however, the field rates of change are sufficiently low for this error to stay small.
Comparing with the results from a detailed multifilamentary strand model in Fig. 3, one can see that a single CS cell is still not sufficiently accurate. Several of them have to be combined to achieve a better matching with the reference, as will be discussed in Section 3.5.
3.4 Inclusion of $j_{\text{c}}(b)$ field dependence
In reality, the magnetization of superconductors decreases with increasing field, as a consequence of the decreasing critical current density $j_{\text{c}}$. Such an effect is not yet reproduced with the hysteresis cells defined above, because the irreversibility parameters $\kappa$ and $\chi$ are still constant values. By making the irreversibility parameters $\kappa$ and $\chi$ decrease with increasing field, one can produce more realistic magnetization curves.
In general, we therefore define field-dependent irreversibility parameters:
$\kappa(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa},\quad\chi(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi},$ (22)
with $\bar{\kappa}$ and $\bar{\chi}$ two constants (in A/m) and $f_{\kappa}$ and $f_{\chi}$ two scaling functions, smoothly evolving from $1$ at zero field to 00 at large fields. The expressions of $f_{\kappa}$ and $f_{\chi}$ are functions of the geometry and must be determined at the parameter identification step. This will be discussed in Section 5.
Note that the introduction of field-dependent irreversibility parameters in the cells brings an additional nonlinearity to the equations to be addressed during the numerical simulation.
@@ 3.5 Chain of cells
The S cell and the CS cell contain the necessary ingredients to reproduce the different regimes observed in the multifilamentary strand, but a single cell alone does not provides a faithful description over a wide field range. Indeed, if the applied magnetic field is smaller than $\kappa$, both the S cell and the CS cell produce exactly zero magnetic flux density, and hence no loss. In reality, the evolution towards saturation is smooth and progressive, and not subject to a single threshold $\kappa$, as in Figs. 10 and 12.
To better approach the smooth magnetization curves of Figs. 4 and 5, we can combine several cells into what we call a chain of cells. To this end, we follow the approach proposed in [27].
The approach consists of decomposing the total magnetic flux density $\boldsymbol{b}$ into a number of $N\in\mathbb{N}_{0}$ contributions $\alpha_{k}\boldsymbol{b}_{k}$, with $k=1,\dots,N$. The $\boldsymbol{b}_{k}$ are described by distinct hysteresis cells, with distinct parameter values and the $\alpha_{k}$ are weights associated with each of them. All cells are subject to the same total magnetic field $\boldsymbol{h}$ and contribute to the total magnetic flux density with weight $\alpha_{k}$. We define
@@
where $\boldsymbol{h}_{\text{rev},k}$ is the reversible field associated with cell $k$. With the mechanical analogy, this corresponds to connecting cells as a chain, i.e., in series, as illustrated in Fig. 14. The power loss is computed cell by cell, in terms of the fractions $\alpha_{k}\dot{\boldsymbol{b}}_{k}$ of the total rate of change $\dot{\boldsymbol{b}}$.
Figure 14: Chain of cells. The total magnetic flux density is computed as $\boldsymbol{b}=\sum_{k=1}^{N}\alpha_{k}\boldsymbol{b}_{k}$. Arbitrary combinations of cells can be put in series. The represented arrangement is relevant for a multifilamentary strand, as explained in Section 7.
The advantage of this approach is that each cell can be solved independently, as they are all subject to the same magnetic field $\boldsymbol{h}$. Fixed point iterations can be performed to account for field-dependent parameters $\kappa$ or $\chi$ that should depend on the total magnetic flux density $\boldsymbol{b}$. The equations for each cell are identical to what was described in Sections 3.2 and 3.3, but each cell only contributes partly to the total magnetic flux density $\boldsymbol{b}$, with a fraction $\alpha_{k}\boldsymbol{b}_{k}$.
Once parameters are identified (see Section 5), the chain of cells defines the hysteresis model, i.e., the constitutive relationship $\boldsymbol{b}=\mathcal{B}(\boldsymbol{h})$ that reproduces the composite strand magnetization and loss.
4 Implementation and Inclusion in FE Framework
The hysteresis model output the magnetic flux density $\boldsymbol{b}(t)$ as a function of a given magnetic field $\boldsymbol{h}(t)$ (from which we can compute the magnetization $\boldsymbol{m}=\boldsymbol{b}/\mu_{0}-\boldsymbol{h}$ and the power loss). Both $\boldsymbol{b}(t)$ and $\boldsymbol{h}(t)$ are vector quantities varying in time. For a numerical simulation, time is discretized and the solutions are sought at successive time steps, based on the knowledge of the solution at previous time steps and on the new value of the magnetic field.
In this section, we first propose update rules for the two types of cells defined in the previous section, the S cell and CS cell. We then discuss the case of a chain of cells and conclude by describing how to include the hysteresis model in a FE framework.

### v2

@@ 3.1 Homogenized model and internal field
Figure 5: Homogenization concept without transport current.
@@
Similarly, the power loss is no longer computed as a Joule loss as in Eq. (2), but is now linked with the magnetic power density $p_{\text{tot}}=\boldsymbol{h}\cdot\dot{\boldsymbol{b}}$. Part of this power is related to stored magnetic energy, and hence is not associated with loss. The complementary part is related to dissipated energy. One major advantage of the energy-based hysteresis model used in this work is that it gives access to both parts (stored and dissipated) separately at all time [27].
@@
where $\boldsymbol{h}_{\text{in}}$ has no immediate meaning in the detailed strand model per se, but is the local magnetic field that the ROHM model will see inside the strand. These curves are shown in Fig. 6. At high frequencies ($f\gtrsim 10^{4}$ Hz), they approach the shape of an ellipse with a slope of $-1$. The area of the closed loops is preserved between Fig. 4 and Fig. 6, as shown in B.
Figure 6: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T. The results are the same as those in Fig. 4 but in terms of $\mu_{0}h_{\text{in}}$ instead of $\mu_{0}h_{\text{app}}$ (see Eq. (7)).
@@ 3.2 Superconductor cell - S cell
As a first step, we present a rate-independent hysteresis cell [27]: the superconductor cell, or S cell. In this cell, the magnetic field $\boldsymbol{h}$ is decomposed into a reversible field, $\boldsymbol{h}_{\text{rev}}$, and an irreversible field [^2] The irreversible field $\boldsymbol{h}_{\text{irr}}$ has nothing to do with the irreversibility field of a superconductor., $\boldsymbol{h}_{\text{irr}}$:
@@
The reversible field defines the magnetic flux density
@@
The irreversible field creates hysteresis by introducing history dependence. Its amplitude $\|\boldsymbol{h}_{\text{irr}}\|$ is bounded by a value $\kappa$ (A/m), the irreversibility parameter. From a known magnetic field $\boldsymbol{h}$, and a given reversible field $\boldsymbol{h}_{\text{rev}}$ that has been established previously (as a function of the history of $\boldsymbol{h}$), the irreversible field $\boldsymbol{h}_{\text{irr}}$ is determined as follows:
@@
where the dot notation (as in $\dot{\boldsymbol{b}}$) represents the time derivative. An illustration of this equation is given in Fig. 7. While the driving field $\boldsymbol{h}$ evolves inside the sphere of radius $\kappa$ centered at $\boldsymbol{h}_{\text{rev}}$, the reversible field $\boldsymbol{h}_{\text{rev}}$ stays constant (and so does $\boldsymbol{b}=\mu_{0}\boldsymbol{h}_{\text{rev}}$), and the irreversible field is defined accordingly, see Fig. 7(a). If $\boldsymbol{h}$ reaches the surface of the sphere, then, the reversible field $\boldsymbol{h}_{\text{rev}}$ (and, hence, also $\boldsymbol{b}$) must evolve in order to maintain the condition $\|\boldsymbol{h}_{\text{irr}}\|\leq\kappa$ valid, and $\boldsymbol{h}_{\text{irr}}$ is established parallel to $\dot{\boldsymbol{b}}$, see Fig. 7(b).
@@
Figure 7: S cell, adapted from [27, 29]. Illustration in 2D of the magnetic field decomposition into reversible and irreversible fields. The dotted elements in (b) refer to the situation at a previous time.
The total power density $p_{\text{tot}}$ (W/m3) is expressed as
@@
The reversible power $p_{\text{rev}}$ is the time derivative of the stored magnetic energy density,
@@
while the irreversible power $p_{\text{irr}}$ is always non-negative and is associated with the dissipated energy density,
@@
A mechanical analogy of the S cell in a one-dimensional setting is shown in Fig. 8. It represents a mechanical system made of a dry friction element in parallel with a restoring spring [27].
Figure 8: Mechanical analogy [27] of the S cell: a parallel connection of a dry friction element and a restoring spring. In the analogy, the system is subject to a force $h$ which causes an elongation $b=\mu_{0}h_{\text{rev}}$.
In practice, the magnetization of superconductors decreases with increasing field as a consequence of the decreasing critical current density $j_{\text{c}}$, as is shown in Fig. 2. In order to reproduce this effect with the S cell, one can make the irreversibility parameter $\kappa$ a function of the magnetic flux density $\boldsymbol{b}$, by defining
$\kappa(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa},$ (14)
with $\bar{\kappa}$ (A/m) a constant and $f_{\kappa}$ a scaling function, smoothly evolving from $1$ at zero field to 00 at large fields. The expression of $f_{\kappa}$ is function of the geometry and must be determined at the parameter identification step, as discussed in Section 5. The introduction of a field-dependent irreversibility parameter brings an additional nonlinearity in the equations, which has to be addressed during the numerical simulation.
Magnetization loops obtained from S cells with different irreversibility parameters are shown Fig. 9 for a unidirectional excitation. By construction, these curves are independent of the rate of change of the applied field.
Figure 9: Hysteresis curves obtained with one S cell with field-dependent or field-independent irreversibility parameters (the function $f_{\kappa}(b)$ is the same as in Section 6). The legend is the same for both subfigures. (Up) Magnetic flux density $b$ versus applied field $\mu_{0}h$. (Down) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$. Note that in this simple case, $h_{\text{irr}}=-m$.
A careful choice of the scaling function $f_{\kappa}(b)$ allows to reproduce the shape of the lower and upper branches of the major hysteresis loop of Fig. 2. However, with a single S cell, the transitions between these upper and lower branches are not smooth as in the reference solution in Fig. 2, but sharp. Also, for applied fields smaller than $\kappa(0)$, the magnetic flux density remains exactly zero and no loss is generated by the S cell. To obtain smooth transitions and better description at small field amplitudes, combining several cells is necessary. This is discussed in Section 3.4.
@@ 3.3 Composite superconductor cell - CS cell
For regimes in which the superconducting strand response is strongly rate-independent, the S cell presented in the previous section must be generalized. Starting from the S cell, we propose to account for the magnetization due to eddy currents and coupling currents by adding two new contributions in the magnetic field decomposition, $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$, such that Eq. (8) is rewritten
@@
Figure 10: CS cell generalizing the S cell by accounting for eddy currents and coupling current effects, as well as for filament coupling effect. Illustration of the magnetic field decomposition into four contributions ($\boldsymbol{h}_{\text{rev}}$, $\boldsymbol{h}_{\text{irr}}$, $\boldsymbol{h}_{\text{eddy}}$, and $\boldsymbol{h}_{\text{coupling}}$). The dashed circle of radius $\chi$ limits the amplitude of the field $\boldsymbol{h}_{\text{coupling}}$. The dotted arrow represents $\boldsymbol{h}_{\text{rev}}$ at a previous time.
Adapting [28], the field $\boldsymbol{h}_{\text{eddy}}$ is defined by
with an eddy current time constant parameter $\tau_{\text{e}}$ (s). The eddy component produces a new loss contribution, $p_{\text{eddy}}$ (W/m3), which reads
In the mechanical analogy, the inclusion of eddy current effects corresponds to adding a dashpot (viscous friction element) in parallel with the dry friction element and the restoring spring, as illustrated in Fig. 10(a). For slow variations of the force $h$, the dashpot is barely opposing the elongation $b$, whereas for faster variations of $h$, it becomes gradually more blocking and starts to oppose variations of $b$. This reproduces magnetic shielding and increased magnetization due to eddy currents.
The other new contribution $\boldsymbol{h}_{\text{coupling}}$ is associated with coupling currents [5]. The coupling currents in the conducting matrix are due to net currents flowing in the superconducting filaments, and are therefore also subject to hysteresis and saturation of the filaments. The field $\boldsymbol{h}_{\text{coupling}}$ has then its own irreversibility parameter $\chi$ (A/m). We define
$\boldsymbol{h}_{\text{coupling}}$ =$&\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}},&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|<\chi,\\ &\chi\ \dot{\boldsymbol{h}}_{\text{rev}}/\|\dot{\boldsymbol{h}}_{\text{rev}}\|,&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|\geq\chi,\\ \end{aligned}\right.$ (19)
with a coupling current time constant $\tau_{\text{c}}$ (s). In order to reproduce reduced magnetization at higher fields, the irreversibility parameter $\chi$ must also be a function of $\boldsymbol{b}$. We therefore define
$\chi(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi},$ (20)
with $\bar{\chi}$ a constant (A/m) and $f_{\chi}(\boldsymbol{b})$ a scaling function. As for $f_{\kappa}(\boldsymbol{b})$, the expression of $f_{\chi}(\boldsymbol{b})$ must be determined in the parameter identification step.
The field $\boldsymbol{h}_{\text{coupling}}$ introduces two new loss contributions, $p_{\text{coupling}}$ and $p_{\text{irr,c}}$ (W/m3), that read
$p_{\text{irr,c}}$ $=\boldsymbol{h}_{\text{coupling}}\cdot\left(\dot{\boldsymbol{b}}-\dot{\boldsymbol{b}}_{\text{coupling}}\right),$ (22)
with
$\dot{\boldsymbol{b}}_{\text{coupling}}=\frac{\mu_{0}}{\tau_{\text{c}}}\boldsymbol{h}_{\text{coupling}}.$ (23)
The mechanical analogy associated with the coupling current contribution is represented in Fig. 10(a). The combination of the dashpot in series with the dry friction element reproduces the saturation of the coupling currents. When the force on the dashpot remains below $\chi$, the dry friction element is at rest and produces no hysteresis. This corresponds to a situation in which the coupling currents are small enough for the hysteresis to be neglected. If the force exceeds $\chi$ on the dashpot, it is then subject to hysteresis. The second hysteresis element, with $\chi$, always acts in parallel with the first one, related to the irreversible field $\boldsymbol{h}_{\text{irr}}$. The magnetization of coupled filaments therefore depends on the sum of both, and so is the associated hysteresis loss. We therefore group $p_{\text{irr}}$ and $p_{\text{irr,c}}$ into a single component $p_{\text{hyst}}=p_{\text{irr}}+p_{\text{irr,c}}$.
@@
In summary, the power loss decomposition reads:
$p_{\text{tot}}=p_{\text{rev}}+\underbrace{p_{\text{irr}}+p_{\text{irr,c}}}_{p_{\text{hyst}}}+p_{\text{eddy}}+p_{\text{coupling}},$ (24)
where $p_{\text{rev}}$ corresponds to stored energy, while the other components correspond to dissipated energy.
Magnetization curves obtained with one CS cell and the associated energy loss per cycle and per unit volume are shown in Figs. 11 and 12. For a given frequency $f$, the energy loss per cycle and per unit volume is computed as
$q_{\text{tot}}=\int_{1/f}^{2/f}p_{\text{tot}}(t)\,\text{d}t.$ (25)
The total loss $q_{\text{tot}}$ is also decomposed in three contributions for easier interpretation, following Eq. (24).
The hysteresis loss exhibits the expected evolution with frequency: two plateaus at low and middle frequencies, with a transition where the coupling loss peaks, and then a sharp decrease when the eddy loss becomes dominant. Coupling and eddy loss curves follow the shape of two bell curves.
Figure 11: Hysteresis curves obtained with one CS cell at different frequencies. Cell parameters are: $\mu_{0}\bar{\kappa}=0.2$ T, $\mu_{0}\bar{\chi}=0.3$ T, $\tau_{\text{c}}=0.2$ s, and $\tau_{\text{e}}=0.3$ ms, and the same functions $f_{\kappa}(b)$ and $f_{\chi}(b)$ as in Section 7.
[…]
In reality, the eddy current loss is asymptotically decreasing as $1/\sqrt{f}$ at high frequencies due to the skin effect. With the CS cell, one can however only expect the curve to decrease as $1/f$ at high frequencies, as a result of Eq. (17). For $f\to\infty$, the CS cell is therefore not expected to provide perfect results and shall be adapted if one wants better results for very high frequencies ($f\gtrsim 10^{3}$ Hz).
For the same reasons as for the S cell, several CS cells must be combined to obtain a good description on a range of field amplitudes. This is discussed in the next section.
@@ 3.4 Chain of cells
The S cell and the CS cell contain the necessary ingredients to reproduce the different regimes observed in a multifilamentary strand, but a single cell alone does not produce a valid response over a wide field range. Indeed, in reality, the evolution towards saturation is smooth and progressive, and not subject to a single threshold field $\kappa$ as is the case in Figs. 9 and 11.
To better approach the smooth magnetization curves of Figs. 2 and 4, we can combine several cells into a chain of cells. To this end, we follow the approach proposed in [27].
The approach consists in decomposing the total magnetic flux density $\boldsymbol{b}$ into $N$ contributions $\alpha_{k}\boldsymbol{b}_{k}$, with $k=1,\dots,N$. Each $\boldsymbol{b}_{k}$ is described by a distinct hysteresis cell, but all cells are driven by the same magnetic field $\boldsymbol{h}$. Each cell contributes to the total magnetic flux density with a weight $\alpha_{k}$. We define
@@
where $\boldsymbol{h}_{\text{rev},k}$ is the reversible field associated with cell $k$. In the mechanical analogy, this corresponds to connecting cells as a chain, i.e., in series, as illustrated in Fig. 13.
Figure 13: (Up) Chain of S cells. (Down) Chain of CS cells.
The power loss is computed by adding contributions of each individual cell, involving the fractions $\alpha_{k}\dot{\boldsymbol{b}}_{k}$ of the total rate of change $\dot{\boldsymbol{b}}$.
The advantage of this approach is that each cell can be solved independently, as they are all subject to the same magnetic field $\boldsymbol{h}$. Fixed point iterations can be performed to account for field-dependent parameters $\kappa$ or $\chi$ that depend on the total magnetic flux density $\boldsymbol{b}$. The equations for each cell are identical to those in Sections 3.2 and 3.3, with the only difference that each cell only contributes partly to the magnetic flux density $\boldsymbol{b}$, with a fraction $\alpha_{k}\boldsymbol{b}_{k}$.
[…]
4 Implementation and Inclusion in a FE Framework
The ROHM model gives the magnetic flux density $\boldsymbol{b}(t)$ as a function of a given magnetic field $\boldsymbol{h}(t)$, from which we can compute the magnetization $\boldsymbol{m}(t)=\boldsymbol{b}(t)/\mu_{0}-\boldsymbol{h}(t)$ and the power loss. All are vector quantities varying in time. For a numerical simulation, time is discretized and the solutions are computed at successive time steps, based on the knowledge of the solution at previous time steps and on the new value of the magnetic field.
In this section, we show how to solve the two types of cells defined in the previous section, the S cell and CS cell, as well as chains of cells. We then describe how to include the ROHM model in a FE framework.